OpenStudy (anonymous):

Find the exact circumference of a circle with an area equal to 36 sq. in. a. 12 pi b. 18 pi c. 324 pi

OpenStudy (anonymous):

12 Pi

OpenStudy (anonymous):

How do you find the radius from this info? Or lack thereof

jimthompson5910 (jim_thompson5910):

A = pi*r^2 36 = 3.14*r^2 solve for r to get r = ??? once you know the value of r, you can find the circumference C C = 2*pi*r

OpenStudy (anonymous):

But how do you find R? Thats where Im stuck

OpenStudy (anonymous):

the question wants the circumference function in pi... so we can get the relation

jimthompson5910 (jim_thompson5910):

did you divide both sides of 36 = 3.14*r^2 by 3.14 ?

OpenStudy (anonymous):

No.

jimthompson5910 (jim_thompson5910):

what do you get when you do so

OpenStudy (anonymous):

Well 36/3.14 is 11.46

jimthompson5910 (jim_thompson5910):

now if r^2 = 11.46, then what must r be?

OpenStudy (anonymous):

What does ^ mean?

jimthompson5910 (jim_thompson5910):

r^2 means r squared

jimthompson5910 (jim_thompson5910):

r^2 = r*r

jimthompson5910 (jim_thompson5910):

for example, 5^2 = 5*5 = 25

OpenStudy (anonymous):

So multiply 11.46 by 11.46?

OpenStudy (anonymous):

Wait, no. 5.73 multiplies by 5.73?

OpenStudy (anonymous):

Multiplied?

jimthompson5910 (jim_thompson5910):

some number, multiplied by itself, gives 11.46

jimthompson5910 (jim_thompson5910):

to find this number, you take the square root of 11.46

jimthompson5910 (jim_thompson5910):

it turns out that sqrt(11.46) = 3.385 roughly which means 3.385*3.385 = 11.458225, pretty close to 11.46

OpenStudy (anonymous):

So then i have 36=3.14*3.38^2?

jimthompson5910 (jim_thompson5910):

now use that r value to find the circumference C use C = 2*pi*r

OpenStudy (anonymous):

So two times pi times 3.38?

jimthompson5910 (jim_thompson5910):

2 times pi times 3.38

jimthompson5910 (jim_thompson5910):

oh sry didn't see your two, nvm

jimthompson5910 (jim_thompson5910):

yes you are correct

OpenStudy (anonymous):

21.23 is what I got.

jimthompson5910 (jim_thompson5910):

Question: does the original "36" given in the problem also have a "pi" attached to it?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok so that changes things a bit

OpenStudy (anonymous):

Sorry.

jimthompson5910 (jim_thompson5910):

if the area is A = 36pi, then A = pi*r^2 36pi = pi*r^2 36 = r^2 ... the pi terms cancel if r^2 = 36, then r = ??

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

yes, then you plug that into C = 2*pi*r

OpenStudy (anonymous):

37.69?

jimthompson5910 (jim_thompson5910):

notice how each answer has a 'pi' in it so ignore the pi for now compute 2*r and then tack on 'pi' at the end

jimthompson5910 (jim_thompson5910):

example: if r = 7, then 2*pi*r = 14pi

OpenStudy (anonymous):

Im still getting the same thing.

jimthompson5910 (jim_thompson5910):

you should get 12pi

OpenStudy (anonymous):

2 times R which is 6 means 12. Then 12*pi. The answer I'm getting is still 37.69

jimthompson5910 (jim_thompson5910):

2*r = 2*6 = 12

jimthompson5910 (jim_thompson5910):

you leave it as 12pi

OpenStudy (anonymous):

OH! Okay.

OpenStudy (anonymous):

Thank you so much! Sorry I am so bad at this!

jimthompson5910 (jim_thompson5910):

whenever you see the key term "exact" that means to leave the 'pi's as they are

jimthompson5910 (jim_thompson5910):

you did good

OpenStudy (anonymous):

Oh. Sounds good. Thank you that means a lot!

jimthompson5910 (jim_thompson5910):

you're welcome